OpenStudy (anonymous):

ok this one is weird..:P Which expression simplifies to 6hthe square root of the quantity 2 times h.?

OpenStudy (gabylovesyou):

what?????????? lol

OpenStudy (anonymous):

oops here \[\sqrt{2h}\]

OpenStudy (anonymous):

simplify

OpenStudy (gabylovesyou):

ohh lol

OpenStudy (anonymous):

lol yea

OpenStudy (gabylovesyou):

h = 0

OpenStudy (gabylovesyou):

exact result is \[ \sqrt{2} \sqrt{h}\]

OpenStudy (anonymous):

um? they are dumb then..:P

OpenStudy (gabylovesyou):

what are your choices lol

OpenStudy (anonymous):

here are the choices \[\sqrt{14h ^{2}} \sqrt{72h ^{3}} \sqrt{36h ^{3}} \sqrt{24h ^{2}}\]

OpenStudy (gabylovesyou):

hold on.. im going to try again lol

OpenStudy (anonymous):

lol okay

OpenStudy (gabylovesyou):

second option lol

OpenStudy (gabylovesyou):

72 h^3

OpenStudy (anonymous):

thank you <3

OpenStudy (gabylovesyou):

np :)